# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ Abstract
Taylor’s hypothesis is the backbone to convert observations done over time to spatial information of the flow while carrying out turbulence measurements on a micrometeorological tower. To address its validity over a highly heterogeneous forest clearcut surface, we utilize an extensive Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) datasets. The DTS measured space-time correlation curves of temperature fluctuations are used to compute the bulk convective speeds ($U_{C}$) of temperature structures in buoyant conditions at a height of $\approx$ 3.1 m above the clearing. These convective speeds are compared with the mean wind speed ($U$) and turbulent intensities of streamwise velocities ($\sigma_{u}\textfractionsolidus U$) obtained from the EC system at the middle of the clearcut. Depending on if $U$ is parallel or perpendicular to the forest edge, the relationships between $U_{C}\textfractionsolidus U$ and $\sigma_{u}\textfractionsolidus U$ are significantly different. However, irrespective of the wind direction, the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion with increasing wavenumbers. The exponent of this power-law ($-0.8$) differs from a homogeneous atmospheric surface layer flow ($-1/3$), thereby pointing towards the effects of heterogeneity. The scale-dependent convective speeds non-linearly transform the temporal frequencies to streamwise wavenumbers, which, eventually, impacts the properties of turbulence (co) spectra. More importantly, this non-linear transformation yields a critical frequency limit, beyond which the eddy length scales derived from frequencies are smaller than the physical dimension of the sonic anemometers, and therefore, cannot be faithfully resolved. This critical limit questions the EC flux estimates beyond this frequency.
@@
The two major components of the Earth’s surface energy budget are sensible and latent heat fluxes. Sans spatial information, the accuracy of these flux measurements is tied to Taylor’s hypothesis, which states that the space-time conversion of any turbulence dataset can be accomplished through the local mean wind speed. We test Taylor’s hypothesis over a highly heterogeneous forest clearcut surface by using fiber-optic cables that acquired air temperatures at high spatial and temporal resolutions. From this data, we estimate the speeds of turbulent eddies at each spatial scale of the flow and derive a critical frequency limit beyond which the space-time conversion is highly uncertain. This critical limit questions the turbulent flux estimates beyond this frequency.
@@ 1 Introduction
The quantification of vertical turbulent transport of heat, matter (e.g. moisture, $\rm CO_{2}$) and momentum are important for understanding the land-atmosphere interaction, energy budget of the Earth’s surface, and carbon and water exchanges between the ecosystem and atmosphere. Under the assumptions of homogeneity and stationary conditions, these vertical transports can be interpreted as surface fluxes, which eventually serve as boundary conditions for the Earth System Models. Given such various applications, the micrometeorological towers at different locations across the globe routinely monitor these fluxes under the aegis of various programs such as FLUXNET, AmeriFlux, NEON, and ICOS (Baldocchi et al., 2001; Novick et al., 2018; Metzger et al., 2019; Heiskanen et al., 2022). Although the turbulent eddies facilitating the vertical transport exist physically in space, spatial information of atmospheric turbulent flows are rarely available and thus the instruments mounted on micrometeorological towers register the turbulent variables over time (Hilland and Christen, 2024). To circumvent this issue, Taylor’s frozen turbulence hypothesis is used to convert the temporal information to spatial ones. This hypothesis states that the turbulent structures of any spatial dimension move past the measurement location with a convection velocity in a frozen state, i.e. their characteristics change slowly compared to the speed at which the mean flow transports them along. In the absence of spatial information, it remains experimentally challenging to verify the accuracy of the Taylor’s hypothesis (TH) under various land-surface conditions.
@@
Despite so, a handful of previous studies attempted to verify TH either through theoretical formulations or by conducting novel field experiments that provided spatial information. From the theoretical side, the early study of Wyngaard and Clifford (1977) used a model developed by Lumley (1965) to investigate the effects of TH on the turbulence statistics at smaller scales of the flow. In particular, Wyngaard and Clifford (1977) provided theoretical formulations that accounted for the deviation from the assumption of constant convective speeds across all scales of the flow. Later on, Tong (1996) experimentally verified their results by using two spatially separated ultrasonic anemometers. Except a few experimental set-ups, spatially explicit sonic anemometer measurements are not readily available (Han et al., 2019). Some studies deployed thermocouples to test TH but their spatial resolution was rather limited (Deardorff and Willis, 1982; Everard et al., 2021). However, with the advent of novel instrumentation, spatial data of turbulent variables can be obtained at high enough resolution either from remote sensing or through fiber optic cables. In that regard, Higgins et al. (2012) used a Raman Lidar to investigate the convective speeds of temperature and moisture structures over a landscape surrounding a water body. On the other hand, Cheng et al. (2017) used distributed temperature sensing (DTS) measurements over a grassland to verify the hypothesis for different atmospheric stability conditions.
@@
The DTS measurement technique uses fiber optic cables and relies on Raman scattering principle to convert the back-scattered energy from the emitted light pulses to temperature values at specific locations in space (Peltola et al., 2021). By laying out these cables either horizontally or vertically, it is therefore possible to obtain temperature data at a fine enough spatial (of the order of tens of cm) and temporal resolutions (of the order of seconds). Cheng et al. (2017) used this spatial temperature data from a horizontal configuration to evaluate scale-dependent convective speeds of turbulent structures. For their experimental dataset over a flat grass surface, they found that at smaller scales of the flow, the turbulent structures moved at speeds that scaled with the streamwise wavenumbers ($\kappa$) as ${(\kappa)}^{-1/3}$. To reach this conclusion, Cheng et al. (2017) adopted the technique of Del Alamo and Jiménez (2009), where differences in Fourier phases of spatial temperature signals at two subsequent instant of time were used to estimate the wavenumber-dependent convection speeds. Cheng et al. (2017) also provided theoretical support to this $-1/3$ exponent by borrowing formulations from the literature of homogeneous and isotropic turbulence. They hypothesized that these scale-dependent convective speeds could enhance the turbulent fluxes by 10-20% at smaller scales of the flow, which, eventually, could solve the crisis of surface energy imbalance.
@@
In recent times, DTS measurements were used over a flat and homogeneous Namib desert surface to verify TH (Hilland and Christen, 2024). Hilland and Christen (2024) found that the TH was almost never satisfied in atmospheric surface layer flows under various stability conditions (ranging from buoyancy- to shear-dominated flows), and the measurement height was an important parameter in that regard. Contrary to Cheng et al. (2017), Hilland and Christen (2024) used a correlation-cube methodology to estimate the scale-dependent convective speeds. In their method, the definition of scales is not related to streamwise wavenumbers in a Fourier domain. In spite of this, Hilland and Christen (2024) found that the convective speeds of smaller-scale turbulent structures were indeed lower than the bulk convective speeds of larger-scale structures. Be that as it may, Hilland and Christen (2024) did not report any power-law variations in their estimates of scale-dependent convective speeds.
@@
All these studies described above were conducted over homogeneous atmospheric surface layer flows. However, in reality, one mostly encounter strong heterogeneity in land-surface conditions, which presents a significant challenge to the atmospheric community (Bou-Zeid et al., 2020). One of such examples is a forest clearcut. Clearcutting is a widely utilized forest management practice to collect timber and in some cases also to restrict the spread of wildfires by limiting the access to the fuel (Ortega et al., 2024). From a fluid dynamical perspective, clearcutting introduces additional complexities to the air flows. For instance, when the air blows directly from the forest canopy sub-layers to clearcut location, it encounters a significant change in surface roughness, which causes the flow to deviate from being in equilibrium with local surface characteristics. The laboratory experiments of Chung and Koseff (2021) over an artificial heterogeneous canopy showed that these deviations are often caused by the canopy-scale eddies extending their footprints on the clearcut surface depending on the distance from the canopy edge. On the other hand, the field-experimental results of Detto et al. (2008) provided evidence that the flow near the forest edge shares several attributes with a backward-facing step flow in which a recirculation zone dominates the turbulence statistics.
@@
From a climate science standpoint, the forests are strong sources and sinks of water vapor and carbon dioxide, respectively, thereby influencing the water- and carbon-cycles between the atmosphere and the ecosystem. Forest clearcutting clearly impacts these exchange processes and the climatic impact of such forest management practices can be observed with the eddy covariance (EC) technique (Tikkasalo et al., 2025). However, due to the aforementioned complexities, the turbulent behavior of the air flow in forest clearcuts pose challenges while interpreting the observed flux values. One of such aspects is regarding the applicability of TH. For instance, if the convective speeds of turbulent structures are indeed scale-dependent then it can have serious consequences on the EC flux measurements as hypothesized by Cheng et al. (2017). Furthermore, the relation derived by Cheng et al. (2017) to describe the scale-dependent convective speeds was under the assumption of homogeneous and isotropic turbulence. A recent study by Chowdhuri and Banerjee (2024) revealed that the turbulence is hardly isotropic at smaller scales of the flow corresponding to the heights closer to the surface. Moreover, it is uncertain whether the assumption of homogeneity would remain valid in highly heterogeneous conditions, such as in a forest clearcut. To address these pertinent issues, our objectives in this study are primarily threefold:
@@
    Estimate the bulk convective speeds of turbulent structures from the DTS data and use that information to validate TH.
@@
    Quantify the impact of scale-dependent convective speeds on turbulence statistics and discuss its implications for EC flux measurements and surface energy balance.
To tackle these objectives, we employ an extensive dataset encompassing both DTS and EC measurements, collected during a five-month field campaign in a forest clearcut. In our analysis, we specifically consider two scenarios depending on whether the direction of the mean wind was perpendicular or parallel to the forest edge. In the latter scenario, when the air blows parallel to the forest edge, it does not encounter sharp changes in surface roughness, and therefore, is expected to behave differently than a forest-edge flow. To estimate the convective speeds of the turbulent structures, the space-time correlation curves are constructed from the DTS temperature measurements. This method has been adopted previously for laboratory (Deardorff and Willis, 1982), large eddy simulation (Su et al., 2000), and field-experimental (Han and Zhang, 2022) datasets. In addition to that, we introduce a novel concept of conditional space-time correlation curves to separately estimate the convective speeds of positive and negative temperature structures. More importantly, we develop a comprehensive framework that accounts for deviations from TH on turbulence spectra and cospectra. This framework, in general, can be applied to situations other than a forest clearcut. The rest of this article is organized in three different sections. In Sect. 2 we describe our measurement set up and methodologies being adopted to verify TH. In Sect. 3, we present and discuss our results, and in Sect. 4 future research directions are outlined.
@@ 2.1 Site description
The measurement campaign took place between May 17 and September 10 2024 at Ränskälänkorpi study site (6111 N, 2516 E, 144 m above sea level), which is located on a peatland that was drained before 1960’s for forestry purposes (Laurila et al., 2021). A Norway spruce (Picea abies) stand growing on the site was clear-felled during late winter - spring of 2021 creating a 6.4 ha clearcut area (see Fig. 1) and Norway spruce seedlings were planted after the clearcutting. Few isolated groupings of living and dead trees were left untouched on the clearcut area and logging debris (e.g. branch piles) were left on site. In general, the terrain of the clearcut is flat, but its surface exhibits prominent small-scale undulations due to e.g. drainage ditches and banks. After clearcutting, pioneer species, such as fireweed (Chamaenerion angustifolium), downy birch (Betula pubescens) and raspberry (Rubus idaeus), have spread on the area and created a varying vegetation cover on the clearing (see Fig. 1a). During the study period, the height of these plants varied approximately between 0.5 m and 1.5 m, depending on location and timing of the growing season. In summary, the clearcut area is a complex mosaic of isolated trees, pioneer species, and logging debris creating a rough surface for the flow.
The clearcut is surrounded by a forest dominated by Norway spruce (Picea abies) with height between 20 m and 25 m depending on the location (see Fig. 1b). Plant area index of the forest in SW direction of the clearing is approximately xx m2 m-2. The forest edges create abrupt changes in canopy height and hence in surface roughness.
@@ 2.2 Measurement setup
The spatial and temporal variability of air temperature above the clearcut was measured with the distributed temperature sensing (DTS) technique Thomas and Selker (2021). The DTS instrument (Ultima-S, 5 km variant, Silixa Ltd., Hertfordshire, UK) was housed in a temperature controlled cabin in the middle of the forest clearcut (see Fig. 1b). The measurements were conducted using double-ended configuration and signals from both ends of the fiber-optic cable (forward and reverse signals) were stored for later processing (Sect. 2.3). The DTS instrument provided observations with 1.2 s temporal and 12.7 cm spatial resolution along the attached cable. Note that in double-ended configuration both ends are measured sequentially and hence after processing the temperature observations were available with 2.3 s resolution.
@@
A white, thin (outer diameter 0.9 mm) aramid reinforced 50 $\mu$m multimode fiber-optic cable (AFL Telecommunications, Duncan, SC 29334, US) was utilized in this study. The cable was arranged in six sections and along each section the temperature values were doubly measured. This was because the measurements were made with a single continuous cable traversing back and forth the different sections (Fig. 1). Sections 1 to 5 were aligned with the forest edge located in South-West direction of the clearcut and the section lengths were between 97 m and 103 m. The cable sections 1, 2, 3, 4 and 5 were approximately at a 20 m, 47 m, 75 m, 115 m and 137 m distance from the forest edge in SW direction, respectively. The sixth section was perpendicular to the other cable sections and to the forest edge and its length was 164 m, out of which 27 m was within the forest. Therefore, the total cable length was 1478 m.
The cable was fastened to wooden support structures and kept at a constant height throughout the measurement domain. The cable locations were determined with a handheld GNSS device (Geo 7X, Trimble Inc., USA). This height was around 3.1 m above the soil, which is similar to the height of the location where eddy covariance (EC) measurements were made (see Fig. 1b). However, due to the heterogeneity in the surface conditions (see Sect. 2.1), local DTS measurement height varied and was approximately between 2 to 4 m above the soil and between 0.5 to 2.5 m above the local roughness elements (e.g. vegetation). Moreover, occasionally the cable was also surrounded by the few isolated trees left at the clearcut.
To quantify the turbulent fluxes over such a complex surface, the EC equipment consisted of a 3-D sonic anemometer (uSonic-3 Cage MP, METEK GmbH, Germany) and an Infrared gas analyzer (LI-7200RS, LI-COR Biosciences, NE, USA). These together measured the three wind velocity components, sonic temperature, and carbon dioxide and water vapor mixing ratios at the same height as DTS, i.e. 3.1 m above the soil. The infrared gas analyzer sampled the flow next to the sonic anemometer and all the EC data were logged with 10-Hz frequency. See more details on the EC measurement setup in Tikkasalo et al. (2025).

### v2

# Taylor’s hypothesis and its impact on turbulence measurements in a forest clearcut

@@ Abstract
Taylor’s hypothesis converts observations done over time to spatial information of the flow while carrying out measurements on a micrometeorological tower. To address its validity over a highly heterogeneous forest clearcut surface, we utilize an extensive Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) datasets. The DTS measured space-time correlation curves of temperature fluctuations are used to compute the bulk convective speeds of temperature structures ($U_{C}$) in buoyant and highly turbulent conditions at a height of $\approx$ 3.1 m above the clearing. These convective speeds are compared with the mean wind speed ($U$) and depending on if $U$ is parallel or perpendicular to the forest edge, the relationships show differences. Particularly, for the parallel-wind scenario, $U_{C}>U$, which we explain by invoking sweeping effects. However, irrespective of the wind direction, the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion (having an exponent of 0.8-0.8) with increasing wavenumbers. These scale-dependent convective speeds impact the properties of temporal turbulence spectrum in the inertial subrange scales by yielding slopes larger than what one would obtain through Taylor’s hypothesis. In our highly turbulent case, these slopes appear enhanced than 5/3-5/3, and we provide a plausible explanation for this by conducting a dimensional analysis. Moreover, the scale-dependent convective speeds non-linearly transform the temporal frequencies to streamwise wavenumbers. This yields a critical frequency limit, beyond which the eddy length scales derived from frequencies are smaller than the physical dimension of the sonic anemometers, and therefore cannot be faithfully resolved.
@@
The measurements of near-surface turbulent flows serve as a cornerstone in land-atmosphere interaction research. Sans spatial information, the accuracy of these turbulence measurements is tied to Taylor’s hypothesis, which states that the space-time conversion can be accomplished through the local mean wind speed. We test Taylor’s hypothesis over a highly heterogeneous forest clearcut surface by using fiber-optic cables that acquired air temperatures at high spatial and temporal resolutions. From this data, we estimate the speeds of turbulent eddies at each spatial scale of the flow and show how that modifies the turbulence statistics when computed over temporal domain. Due to a non-linear transformation of time to space, a critical frequency limit exists beyond which the space-time conversion appears uncertain.
@@ 1 Introduction
The quantification of vertical turbulent transport of heat, matter (e.g. moisture, $\rm CO_{2}$) and momentum are important for understanding the land-atmosphere interaction, energy budget of the Earth’s surface, and carbon and water exchanges between the ecosystem and atmosphere. Under the assumptions of homogeneity and stationary conditions, these vertical transports can be interpreted as surface fluxes, which eventually serve as boundary conditions for the Earth System Models. Given such various applications, the micrometeorological towers at different locations across the globe routinely monitor these fluxes under the aegis of various programs such as FLUXNET, AmeriFlux, NEON, and ICOS (baldocchi2001fluxnet; novick2018ameriflux; metzger2019neon; heiskanen_integrated_2022). Although the turbulent eddies facilitating the vertical transport exist physically in space, spatial information of atmospheric turbulent flows are rarely available and thus the instruments mounted on micrometeorological towers register the turbulent variables over time (hilland2024systematic). To circumvent this issue, Taylor’s frozen turbulence hypothesis is used to convert the temporal information to spatial ones. This hypothesis states that the turbulent structures of any spatial dimension move past the measurement location with a convection velocity in a frozen state, i.e. their characteristics change slowly compared to the speed at which the mean flow transports them along. In the absence of spatial information, it remains experimentally challenging to verify the accuracy of the Taylor’s hypothesis (TH) under various land-surface conditions.
@@
Despite so, a handful of previous studies attempted to verify TH either through theoretical formulations or by conducting novel field experiments that provided spatial information. From the theoretical side, the early study of wyngaard1977taylor used a model developed by lumley1965interpretation to investigate the effects of TH on the turbulence statistics at smaller scales of the flow. In particular, wyngaard1977taylor provided theoretical formulations that accounted for the deviation from the assumption of constant convective speeds across all scales of the flow. Later on, tong1996taylor experimentally verified their results by using two spatially separated ultrasonic anemometers. Except a few experimental set-ups, spatially explicit sonic anemometer measurements are not readily available (han2019applicability). Some studies deployed thermocouples to test TH but their spatial resolution was rather limited (deardorff1982investigation; everard2021sweeping). However, with the advent of novel instrumentation, spatial data of turbulent variables can be obtained at high enough resolution either from remote sensing or through fiber optic cables. In that regard, higgins2012effect used a Raman Lidar to investigate the convective speeds of temperature and moisture structures over a landscape surrounding a water body. On the other hand, cheng2017failure used distributed temperature sensing (DTS) measurements over a grassland to verify the hypothesis for different atmospheric stability conditions.
@@
The DTS measurement technique uses fiber optic cables and relies on Raman scattering principle to convert the back-scattered energy from the emitted light pulses to temperature values at specific locations in space (peltola2021suitability). By laying out these cables either horizontally or vertically, it is therefore possible to obtain temperature data at a fine enough spatial (of the order of tens of cm) and temporal resolutions (of the order of seconds). cheng2017failure used this spatial temperature data from a horizontal configuration to evaluate scale-dependent convective speeds of turbulent structures. For their experimental dataset over a flat grass surface, they found that at smaller scales of the flow, the turbulent structures moved at speeds that scaled with the streamwise wavenumbers ($\kappa$) as ${(\kappa)}^{-1/3}$. To reach this conclusion, cheng2017failure adopted the technique of del2009estimation, where differences in Fourier phases of spatial temperature signals at two subsequent instant of time were used to estimate the wavenumber-dependent convection speeds. cheng2017failure also provided theoretical support to this 1/3-1/3 exponent by borrowing formulations from the literature of homogeneous and isotropic turbulence. They hypothesized that these scale-dependent convective speeds could enhance the turbulent fluxes by 10-20% at smaller scales of the flow, which, eventually, could solve the crisis of surface energy imbalance.
@@
In recent times, DTS measurements were used over a flat and homogeneous Namib desert surface to verify TH (hilland2024systematic). hilland2024systematic found that the TH was almost never satisfied in atmospheric surface layer flows under various stability conditions (ranging from buoyancy- to shear-dominated flows), and the measurement height was an important parameter in that regard. Contrary to cheng2017failure, hilland2024systematic used a correlation-cube methodology to estimate the scale-dependent convective speeds. In their method, the definition of scales is not related to streamwise wavenumbers in a Fourier domain. In spite of this, hilland2024systematic found that the convective speeds of smaller-scale turbulent structures were indeed lower than the bulk convective speeds of larger-scale structures. Be that as it may, hilland2024systematic did not report any power-law variations in their estimates of scale-dependent convective speeds.
@@
All these studies described above were conducted over homogeneous atmospheric surface layer flows. However, in reality, one mostly encounter strong heterogeneity in land-surface conditions, which presents a significant challenge to the atmospheric community (bou2020persistent). One of such examples is a forest clearcut. Clearcutting is a widely utilized forest management practice to collect timber and in some cases also to restrict the spread of wildfires by limiting the access to the fuel (ortega2024modeling). From a fluid dynamical perspective, clearcutting introduces additional complexities to the air flows. For instance, when the air blows directly from the forest canopy sub-layers to clearcut location, it encounters a significant change in surface roughness, which causes the flow to deviate from being in equilibrium with local surface characteristics. The laboratory experiments of chung2021turbulence over an artificial heterogeneous canopy showed that these deviations are often caused by the canopy-scale eddies extending their footprints on the clearcut surface depending on the distance from the canopy edge. On the other hand, the field-experimental results of detto2008structure provided evidence that the flow near the forest edge shares several attributes with a backward-facing step flow in which a recirculation zone dominates the turbulence statistics.
@@
From a climate science standpoint, the forests are strong sources and sinks of water vapor and carbon dioxide, respectively, thereby influencing the water- and carbon-cycles between the atmosphere and the ecosystem. Forest clearcutting clearly impacts these exchange processes and the climatic impact of such forest management practices can be observed with the eddy covariance (EC) technique (tikkasalo2025eddy). However, due to the aforementioned complexities, the turbulent behavior of the air flow in forest clearcuts pose challenges while interpreting the observed flux values. One of such aspects is regarding the applicability of TH. For instance, if the convective speeds of turbulent structures are indeed scale-dependent then it can have serious consequences on the EC flux measurements as hypothesized by cheng2017failure. Furthermore, the relation derived by cheng2017failure to describe the scale-dependent convective speeds was under the assumption of homogeneous and isotropic turbulence. A recent study by chowdhuri2024quantifying revealed that the turbulence is hardly isotropic at smaller scales of the flow corresponding to the heights closer to the surface. Moreover, it is uncertain whether the assumption of homogeneity would remain valid in highly heterogeneous conditions, such as in a forest clearcut.
In many respects, forest clearcut flows share resemblance with roughness sublayer flows (flesch_wind_1999). The study by everard2021sweeping on roughness sublayer flows over a vineyard canopy has suggested that when the turbulence intensities are high, strong sweeping effects associated with large eddies of the flow can cause severe modifications to the convective speeds of smaller-scale eddies. This potentially could cause troubles while interpreting the high-frequency EC measurements of turbulence statistics. To address these pertinent issues, our objectives in this study are primarily threefold:
@@
    Estimate the bulk convective speeds and sweeping speeds of turbulent structures in roughness sublayer flows from the DTS data and use that information to validate TH.
@@
    Quantify the impact of scale-dependent convective speeds on turbulence statistics and discuss its implications for EC measurements.
To tackle these objectives, we employ an extensive dataset encompassing both DTS and EC measurements, collected during a five-month field campaign in a forest clearcut. In our analysis, we specifically consider two scenarios depending on whether the direction of the mean wind was perpendicular or parallel to the forest edge. In the latter scenario, when the air blows parallel to the forest edge, it does not encounter sharp changes in surface roughness, and therefore, is expected to behave differently than a forest-edge flow. To estimate the convective speeds of the turbulent structures, the space-time correlation curves are constructed from the DTS temperature measurements. This method has been adopted previously for laboratory (deardorff1982investigation), large eddy simulation (su2000two), and field-experimental (everard2021sweeping; han2022applicability) datasets. To compute the sweeping speeds, we follow everard2021sweeping. They showed that the sweeping effects can be quantified by employing the elliptic approximation model of he2010small to describe the space-time correlation curves. In fact, everard2021sweeping provided a recipe to compute the sweeping speeds from field-experimental data. Apart from these computations, we develop a practical framework that accounts for deviations from TH on turbulence spectra and cospectra. This framework could be applied to situations other than a forest clearcut. The rest of this article is organized in three different sections. In Sect. 2 we describe our measurement set up and methodologies being adopted to verify TH. In Sect. 3, we present and discuss our results, and in Sect. 4 future research directions are outlined.
@@ 2.1 Site description
The measurement campaign took place between May 17 and September 10 2024 at Ränskälänkorpi study site (6111 N, 2516 E, 144 m above sea level), which is located on a peatland that was drained before 1960’s for forestry purposes (laurila_set-up_2021). A Norway spruce (Picea abies) stand growing on the site was clear-felled during late winter - spring of 2021 creating a 6.4 ha clearcut area (see Fig. 1) and Norway spruce seedlings were planted after the clearcutting. Few isolated groupings of living and dead trees were left untouched on the clearcut area and logging debris (e.g. branch piles) were left on site. In general, the terrain of the clearcut is flat, but its surface exhibits prominent small-scale undulations due to e.g. drainage ditches and banks. After clearcutting, pioneer species, such as fireweed (Chamaenerion angustifolium), downy birch (Betula pubescens) and raspberry (Rubus idaeus), have spread on the area and created a varying vegetation cover on the clearing (see Fig. 1a). During the study period, the height of these plants varied approximately between 0.5 m and 1.5 m, depending on location and timing of the growing season. In summary, the clearcut area is a complex mosaic of isolated trees, pioneer species, and logging debris creating a rough surface for the flow.
The clearcut is surrounded by a forest dominated by Norway spruce (Picea abies) with height between 20 m and 25 m depending on the location (see Fig. 1b). Leaf area index of the forest in SW direction of the clearing is approximately 2.4 m2 m-2. The forest edges create abrupt changes in canopy height and hence in surface roughness.
@@ 2.2 Measurement setup
The spatial and temporal variability of air temperature above the clearcut was measured with the distributed temperature sensing (DTS) technique thomas_optical_2021. The DTS instrument (Ultima-S, 5 km variant, Silixa Ltd., Hertfordshire, UK) was housed in a temperature controlled cabin in the middle of the forest clearcut (see Fig. 1b). The measurements were conducted using double-ended configuration and signals from both ends of the fiber-optic cable (forward and reverse signals) were stored for later processing (see Sect. 22.3). The DTS instrument provided observations with 1.2 s temporal and 12.7 cm spatial resolution along the attached cable. Note that in double-ended configuration both ends are measured sequentially and hence after processing the temperature observations were available with 2.3 s resolution.
@@
[…]
Due to heterogeneity in the surface conditions (see Sect. 22.1), the cable height above the soil changed from one location to another. To avoid any terrain-induced curvature effects on the cable orientation, we decided not to follow the local terrain but to fasten the cable to wooden support structures at a constant height above sea level. This height was identified at each location with a handheld GNSS device (Geo 7X, Trimble Inc., USA), utilized to position the support structures in the area. At the location of the eddy covariance (EC) station, the cable height was around 3.1 m above the soil (see Fig. 1b). However, the local DTS measurement height varied and was approximately between 2 to 4 m above the soil and between 0.5 to 2.5 m above the local roughness elements (e.g. vegetation). Moreover, occasionally the cable was also surrounded by the few isolated trees left at the clearcut. For practical purposes, we ignore these local height variations and consider the height of the DTS cable to be 3.1 m above the soil in order to be consistent with the EC measurements.
To quantify the turbulent fluxes over such a complex surface, the EC equipment consisted of a 3-D sonic anemometer (uSonic-3 Cage MP, METEK GmbH, Germany) and an Infrared gas analyzer (LI-7200RS, LI-COR Biosciences, NE, USA). These together measured the three wind velocity components, sonic temperature, and carbon dioxide and water vapor mixing ratios at the height of 3.1 m above the soil. The infrared gas analyzer sampled the flow next to the sonic anemometer and all the EC data were logged with 10-Hz frequency. See more details on the EC measurement setup and flux footprint calculations in tikkasalo2025eddy.